# From Tokens to Words: On the Inner Lexicon of LLMs

### v2

@@ 4.2 Multi-Token Words
We have seen that LLMs are able to detokenize single-token words that are artificially broken into sub-words into their correct (single-token) word. This indicates that these words are part of the model’s inner lexicon. But what about multi-token words? By definition, these words do not have a single embedding vector in the BPE vocabulary, so we cannot compare them to any existing vector, and thus cannot use logit lens.
Instead, we use the Patchscopes technique (Ghandeharioun et al., 2024), which interprets the hidden representation of a model using the model’s language abilities. In particular, we feed the model with the prompt “Repeat this word twice: 1) $X$ 2)”, where $X$ is the hidden representation of the last word token. Our hypothesis is that if the multi-token word is found in the model internal lexicon, than it will “understand” its representation in the input layer as well, and successfully repeat it by generating all tokens of the original word (see Fig. 1, right).
We evaluate the model by computing the rate of retrieval—the proportion of times it repeats the correct (multi-token) word. We use the same models and dataset as in the previous experiments, though in this case we do not further split or add typos, as the words are already split into multiple tokens. Particularly, we feed each multi-token word (without any context) to the model, and extract the last token’s representation from each layer. We then use Patchscopes’ prompt in a new run to check if the model regenerates the original multi-token word.
Our Llama2-7B results (Fig. 3(b)) show a striking trend: when feeding the model with vectors from layers 5–7, it is able to repeat the word in 64% of the cases, despite never seeing this vector as input before.77777.4% of the words are repeated correctly in any layer. As to the different layers, we observe a similar trend to the single-token results: retrieval rates increase rapidly at some point (here earlier than in previous experiments, around layer 3-4), reach a peak at layer 7, and then start to decline. Combined, these results suggest that the model treats multi-tokens words as if they were in the vocabulary but split into sub-word tokens, indicating a latent vocabulary that expands beyond the tokenizer’s limitations. Interestingly, we observe that 22.6% of the multi-token words are not successfully decoded from any internal layer, hinting that they might not be represented in the model’s inner lexicon.
@@ 5 How Does Detokenization Happen?
We have so far shown that LLMs perform a detokenization step: at some point, the hidden representation of the final token of a given word becomes strikingly similar to the (single-word) vector of that word. This process is robust to artificial splits of single token words, to splits due to typos, and even to multi-token words, which the model can still recognize at the input layer, despite never having seen them there during training.
We next turn to ask how does the model reconstructs full word representations from sub-word tokens? We aim to understand the dynamics of this process by analyzing the main transformer components—feedforward network layers (FFN) and the attention mechanism.
@@ 5.1 Word Retrieval using the Feedforward Mechanism
FFN layers have been shown to serve as key-value memories for storing memory (Geva et al., 2022a; 2021; Burns et al., 2024). We hypothesize that this memory might be used to store the inner lexicon as well. Particularly, we suggest that the model uses the FFNs to refine the representation of the final token, enabling it to retrieve the original word.
To test this hypothesis, we repeat the typos experiment of single-token words (Section 4.1), but this time applying logit lens to the output of the FFN layers, instead of the hidden representation. Figure 4(a) shows the retrieval rate of the original word. We observe that the retrieval of the full word concepts from the FFN layers appears a few layers earlier compared to the hidden state, which indicates that they help refine it. Nonetheless, the rate of retrieval appears to be lower in FFNs, which suggests that other factors might come into play in building the hidden representation of words. However, in Fig. 4(b) we plot the cumulative retrieval rate, i.e., for each layer, whether the word was identified in any layer so far. Our results indicate a different conclusion: the FFN update vectors match the (single-token) input representation of the word 70% of the time in layer, particularly both earlier and more frequently compared to the hidden representation. This suggest a larger variance in the exact layer in which FFNs identify the word compared to the hidden state. Our results indicate that the FFNs play a pivotal role in reconstructing word-level information from sub-word tokens.
5.2 Token Aggregation
Our results hint that LLMs retrieve words form the FFN layers, which allows the final token to represent these words. But this role of the FFN layers only starts to emerge in layers 3–4. What happens earlier? We rely on previous work that showed that in the early layers the model focuses on integrating information from nearby tokens to compose entities (Lad et al., 2024), and hypothesize that the model starts by aggregating information from the previous sub-word tokens.
@@ 5.2 Token Aggregation
Figure 4: Word retrieval from the FFN layers vs. from the hidden state. (4a) Retrieval rates across layers. The FFN values peak before word retrieval begins in the hidden layers. (4b) Its cumulative version, showing that word retrieval occurs earlier and more frequently in FFNs than in the hidden representation.
To test this, we take all 2-token words in a subset of 1,000 documents from WikiText-103 (a total of 5,571 words), and feed them (along with their context) to the Llama2-7B model. We then measure the average attention weights of the final token to the prefix tokens in each layer. As control, we measure the average attention weights assigned by single-token words to their preceding token.
@@
Figure 5: Attention weights for 2-token words: Early peaks (layers 2–3) show high attention values from the 2nd sub-word token to the first, but these values decline rapidly. Attention from single-token words to their previous token shows a similar trend, though with substantially lower values at first, which become higher later.
@@
Our results (Fig. 5) support previous findings (Lad et al., 2024)—the attention to previous tokens is high in the first 2–3 layers, but then declines sharply (up to $\sim$90decline). [^8] Experiments with 3- and 4-token words show a similar trend, see Appendix B. Considering the attention from single-word tokens to their previous token, we also observe this peak phenomenon, but importantly—here the initial peak is far lower than in our multi-token words [^9] The diverging patterns between single and multi-token words are statistically significant; see Appendix B.. However, we also note that the following values are in turn higher. These results suggest that LLM strongly attend to their sub-word tokens at first, but then largely ignore them.
@@
Our results suggest that the model performs a detokenization process by first aggregating information from the prefix tokens into the final token’s hidden representation. It then refines the representation of the final token in the FFN layers to retrieve the full word. This two-stage process of token aggregation and concept retrieval provides insight into the internal mechanisms of LLMs and how they handle sub-word tokens to reconstruct word-level representations.
@@ 6 Expanding LLM Vocabulary without Finetuning
We have shown that language models internally fuse multi-token words into single-token representations, which models can decode to generate the complete word. This raises the question whether models can make more general use of such fused representations when processing textual input, by using them in place of the original, multi-token embeddings. Conversely, models were shown to foreshadow several future tokens in a single hidden state without being explicitly trained to do so (Pal et al., 2023); can we leverage the fused representations to enable models to generate multi-token words in a single inference step?
We next explore whether such vocabulary expansion is possible without any updates to the model’s parameters. [^10] We do add new entries to the model’s embedding and unembedding matrices, but these are inherently required to expand the vocabulary and represent the new vocabulary words. The potential of such methods is high, as even tokenizers intentionally trained for multilingual support result in extremely longer tokenization lengths for non-English languages (up to 13x for some languages), affecting cost and latency (Ahia et al., 2023; Sengupta et al., 2023; Petrov et al., 2024). Still, previous attempts to expand tokenizer vocabulary are limited in number, and require further training on large corpora (Kim et al., 2024; Zhao et al., 2024) or attempt to force alignment between languages in embedding space (Wang et al., 2019), rather than rely on the internal detokenization mechanism of LLMs.
We first experiment with expanding the input embeddings matrix $E$, leaving the output unembedding matrix $U$ unchanged. To extract fused representations, we begin with the same approach as in Section 4.2: Given a multi-token word, we run it through the model using a prompt template $P$ [^11] We use the template “x x x x” where x is the new vocabulary word; preliminary experiments showed it provides better results than the template “x”, possibly due to the effect a token’s position has on its representation. We then identify the first layer in which the last token is successfully decoded into the full word (Ghandeharioun et al., 2024). We extract the hidden state of the last token in that layer, and add it as the word’s fused representation to $E$. If the word is never successfully decoded from the representation in any of the layers, we don’t attempt to add it to the vocabulary, as we assume it is not in the model’s inner lexicon. We experiment with two approaches: identity, where we use the fused representation as is; and linear, where we first transform it with a linear map $T_{\ell,E}$. We learn $T_{\ell,E}$ by fitting a linear regression from the layer $\ell$ hidden states of all single-token words when passed with the template $P$, to their corresponding input embeddings. Importantly, $T_{\ell,E}$ is both not part of the original model, and does not depend on the new, originally multi-token words we add to the vocabulary.
Original vocabulary tokens Newly added words
Top-1 Acc. Mean Reciprocal Rank Top-1 Acc. Mean Reciprocal Rank
Original vocabulary 57.7% 0.679 47.5% 0.556
Last token embedding ($E$) 54.5% 0.647
Fused-Identity ($E$) 53.5% 0.638
Fused-Linear ($E$) 56.7% 0.669
Fused-Linear ($E$ + $U$) 56.7% 0.667 19.5% 0.598
Table 1: Performance of finetuning-free vocabulary expansion for Llama2-7B on WikiText-103. Different rows correspond to baselines (first two rows) and proposed approaches (last three). Methods with $E$ add new multi-token words as single-token entries in the embedding matrix, while those with $U$ add new entries to the unembedding matrix. Column groups correspond to evaluations of language model performance on different token groups: those already in the vocabulary (left) and the newly added entries (right, relevant only when expanding $U$). Results marked with $\ast$ are for predicting the 1st token of a multi-token word (rather than predicting the new entry of the complete word). Our best approach (Fused-Linear) leverages the fused representations emerging within models after detokenization, to add new vocabulary entries for multi-token words. We show it maintains performance on the original vocabulary, and enables the model to successfully utilize entries in $U$ that were never observed during training.
We apply our approach to expand the vocabulary of Llama2-7B with all multi-token words found in the WikiText-103 test set, resulting in 4.3k newly added words. To examine whether the model can correctly use the new entries, we evaluate it in a language modeling setup, when all newly added words are replaced with their fused single-token embeddings, similar to our setup in Section 4.2 (see righthand side in Fig. 1). We compare our approach to two baselines: using the original vocabulary without any new entries; and using the unprocessed last token embedding of each new word as its new vocabulary entry (instead of the fused representations). For evaluation, we cannot use perplexity, the standard LM measure, as the different approaches have different vocabularies, which influences perplexity scores and prevents a fair comparison. Instead, we measure the token-level top-1 accuracy and the mean reciprocal rank of the ground truth tokens; naturally, we do not report the ranking of sub-word tokens forming the newly added words, as they only serve as ground truth in our baseline.
We report our results in Table 1. Examining model performance when using the fused representations in the identity approach, we find that while the model can successfully decode the full word information when explicitly prompted to do so, it fails to utilize the fused representation for inference, and achieves slightly worse performance than when using the unprocessed last token embeddings. However, using the linear approach by first applying $T_{\ell,E}$ transforms the fused representation into a usable form for initial layers, as we find the model can then utilize the newly added words and maintains its performance. [^12] A possible explanation for the importance of $T_{\ell,E}$ is that input embeddings have very different distributions from intermediate hidden states, as LLMs are extremely sensitive to 1st layer ablations (Lad et al., 2024).
Given the model’s ability to use fused representations as inputs, we next enable the model to generate multi-token words in a single inference step, by adding new entries to the unembedding matrix $U$ derived from their fused representations. In light of our previous results, we only experiment with the linear approach. We follow the same steps as before, this time applying linear maps $T_{\ell,U}$ which we learn by fitting linear regressions from the layer $\ell$ hidden states of all existing single-token words, to their corresponding rows in the unembedding matrix. We evaluate whether the model chooses to generate the new vocabulary entries, or rather the first token of the original multi-token words (as it does when using the unedited vocabulary).
Surprisingly, we find the model assigns, on average, higher rank to the new vocabulary entries over their original 1st sub-word tokens in the appropriate contexts (Table 1, last row). While the top-1 accuracy is higher for the 1st sub-word tokens in the unedited vocabulary, we find that when new vocabulary entries are not assigned high ranks, the original sub-word will often be predicted. We further observe that in 7.1% of examples, the new entry achieves the top-1 rank even though the original sub-word in the baseline does not. Notably, this also does not compromise performance on existing vocabulary words. Our results suggest that the model can sometimes try to generate the fused representation, but as it does not exist in its (original) unembedding vocabulary, the first sub-word token is selected instead. Overall, our findings show potential to substantially decrease inference times for LLMs. In future work, we will explore their application to multilingual LLMs (Petrov et al., 2024).

### v3

@@ 4.2 Multi-Token Words
We have seen that LLMs are able to detokenize single-token words that are artificially broken into sub-words back into their correct (single-token) word. This indicates that these words are part of the model’s inner lexicon. But what about multi-token words? By definition, these words do not have a single embedding vector in the BPE vocabulary, so we cannot compare them to any existing vector, and thus cannot use logit lens.
Instead, we use the Patchscopes technique (Ghandeharioun et al., 2024), which interprets the hidden representation of a model using the model’s language abilities. In particular, we feed the model with the prompt “Repeat this word twice: 1) $X$ 2)”, where $X$ is the hidden representation of the last word token. Our hypothesis is that if the multi-token word is found in the model’s internal lexicon, then it will “understand” its representation in the input layer as well, and successfully repeat it by generating all tokens of the original word (see Fig. 1, right).
We evaluate the model by computing the rate of retrieval—the proportion of times it generates the correct (multi-token) word. We use the same models and dataset as in the previous experiments, though in this case we do not further split or add typos, as the words are already split into multiple tokens. Particularly, we feed each multi-token word (along with its 100-tokens context, as before) to the model, and extract the last token’s representation from each layer. We then use Patchscopes’ prompt in a new run to check if the model regenerates the original multi-token word.
Our Llama2-7B results (Fig. 3(b)) show a striking trend: when feeding the model with vectors from layers 5–7, it is able to repeat the word in 64% of the cases, despite never seeing this vector as input before. [^9] Overall, 77.4% of the words are repeated correctly in at least one layer. As to the different layers, we observe a similar trend to the single-token results: retrieval rates increase rapidly at some point (here earlier than in previous experiments, around layer 3-4), reach a peak at layer 7, and then start to decline. Combined, these results suggest that the model treats multi-tokens words as if they were in the vocabulary but split into sub-word tokens, indicating a latent vocabulary that expands beyond the tokenizer’s limitations. Interestingly, we observe that 22.6% of the multi-token words are never successfully decoded from any internal layer, hinting that they might not be represented in the model’s inner lexicon.
@@ 5 How Does Detokenization Happen?
We have so far shown that LLMs perform a detokenization step: at some point, the hidden representation of the final token of a given word becomes strikingly similar to the (single-tokenx) vector of that word. This process is robust to artificial splits of single token words, to splits due to typos, and even to multi-token words, which the model can still recognize at the input layer, despite never having seen them there during training.
We next turn to ask how does the model reconstruct full word representations from sub-word tokens? We aim to understand the dynamics of this process by analyzing the main transformer components: feedforward network layers (FFN) and the attention mechanism.
@@ 5.1 Word Retrieval Using the Feedforward Mechanism
FFN layers have been shown to serve as key-value tables for storing memory (Geva et al., 2022; 2021; Meng et al., 2022b). We hypothesize that this memory might be used to store the inner lexicon as well. Particularly, we suggest that the model uses the FFNs to refine the representation of the final token, enabling it to retrieve the original word.
@@
Figure 4: Word retrieval in FFN layers vs. in the hidden states. (4a) Retrieval rates across layers. The FFN values peak before word retrieval begins in the hidden layers. (4b) Its cumulative version, showing word retrieval occurs earlier and more frequently in FFNs than in the hidden representation.
To test this hypothesis, we repeat the typos experiment of single-token words (Sec. 4.1), but this time applying logit lens to the output of the FFN layers, instead of the hidden representation. Figure 4(a) shows the retrieval rate of the original word. We observe that the retrieval of the full word concepts from the FFN layers occurs a few layers earlier compared to the hidden state, which indicates that they help refine it. Nonetheless, the rate of retrieval appears to be lower in FFNs, which suggests that other factors might come into play in building the hidden representation of words. However, in Fig. 4(b) we plot the cumulative retrieval rate, i.e., for each layer, whether the word is identified in any layer so far. Our results indicate a different conclusion: the FFN update vectors match the (single-token) input representation of the word 70% of the time in any layer, particularly both earlier and more frequently compared to the hidden representation. This hints that FFNs indeed play a substantial role in building the internal word representations in LLMs.
We further investigate the role FFNs play in detokenization through ablation experiments on FFN updates in Llama2-7B. Specifically, we test whether these updates are necessary for the word representation to emerge in the residual stream. To do so, we artificially split single-token words in a similar setting to Sec. 4.1, this time focusing on derived words, formed by adding a suffix to a root word. [^10] We examine words with three common suffixes: “ing”, “ion”, and “est”. We split these back to two parts, for example, we divide “eating” to “eat” and “ing”, so that processing the suffix token is essential to reconstructing the word correctly. Using logit lens, we examine the FFN updates to the suffix token’s residual stream, and selectively ablate those associated with the original single-token word ($\sim$5of layers). As a control, for each word, we ablate an equal number of random FFN updates. Our results (Fig. 12 in App. E) show that removing the updates carrying full-word representations dramatically reduces retrieval rates—from 85% without ablation to just 18%. In contrast, ablating random FFN updates has little to no effect.
This indicates that FFN updates are essential for a detokenized representation to emerge in the final token. But does this affect the model’s ability to ”understand” the word, particularly when predicting the next token? To investigate this, we evaluate the model’s ability to retrieve the capital cities of countries, using the prompt “The capital of [COUNTRY] is ______”. We use single-token country names artificially split into two tokens, and follow the previous ablation protocol. We observe notable patterns: randomly ablating FFN updates reduces performance from 88% to 74%, [^11] We note that baseline performance when passing country names without artificial splits is 95%. likely due to disrupting the model’s factual recall mechanism (Meng et al., 2022b). However, specifically canceling updates detected as country’s name (similarly occurring in only $\sim$5of layers) causes a sharp drop to 41%. Overall, our results suggest that FFNs are central to detokenization, and actively reconstruct full-word meanings in LLMs.
5.2 Token Aggregation
Our results suggest that LLMs use the FFN layers to store and retrieve word representations, which are accessed in the final token to reconstruct the complete word. However, this role of the FFN layers only begins to emerge around layers 3–4. What happens earlier? We build on previous work showing that early layers primarily integrate information from nearby tokens to compose entities (Lad et al., 2024), and hypothesize that the model starts by aggregating information from the previous sub-word tokens.
@@ 5.2 Token Aggregation
Figure 5: Attention weights for 2-token words: Early peaks (layers 2–3) show high attention values from the second sub-word token to the first, but these values decline rapidly. Attention from single-token words to their previous token shows a similar trend, though with substantially lower values at first, which become higher later.
To test this, we extract all two-token words in a subset of 1,000 WikiText-103 documents (a total of 5,571 words), and feed them, along with their context, to Llama2-7B. We then measure the average attention weights of the final token to the prefix token in each layer. As a control, we also measure the average attention weights assigned by single-token words to their preceding token.
@@
Our results (Fig. 5) support previous findings (Lad et al., 2024)—the attention to previous tokens is high in the first 2–3 layers, but then declines sharply (by up to $\sim$90%). [^12] Experiments with 3- and 4-token words show a similar trend, see App. B. For single-token words, we observe a similar attention pattern to the previous token in the first layers, but importantly—the initial peak is significantly lower than in multi-token words. [^13] The diverging patterns between single and multi-token words are statistically significant; see App. B. Still, in later layers, the attention weights of single-token words are in turn higher than for multi-token words. These results suggest LLMs strongly attend to the preceding sub-word tokens of multi-token words at first, but then largely ignore them.
@@
Altogether, our results suggest that LLMs perform a detokenization process by first aggregating information from the prefix tokens into the final token’s hidden representation, and then refining the representation of the final token using the FFN layers to retrieve the full word’s concept representation. This two-stage process of token aggregation and concept retrieval provides insight into the mechanisms LLMs use to handle sub-word tokens and reconstruct word-level representations.
@@ 6 Expanding LLM Vocabulary Without Finetuning
We have shown that language models internally fuse multi-token words into a single-token representation, and that they can further “read” these representations as inputs, and decode the original multi-token words. This raises the question: can models use these fused representations instead of the original multi-token inputs—to encode input prompts using less tokens and reduce computation? Similarly, models were shown to implicitly predict several future tokens in a single hidden state without being explicitly trained to do so (Pal et al., 2023); can we leverage these representations to enable models to predict multi-token words in a single inference step?
Motivated by our findings, we explore whether we can expand the model’s vocabulary with new input and output embeddings for originally multi-token words, without any updates to model parameters. [^14] We do add new entries to the model’s embedding and unembedding matrices, but these are intrinsically required to expand the vocabulary and represent the new vocabulary words. This goal is of practical importance, especially for low-resource languages and domains: even tokenizers built for multilingual support often produce substantially longer token sequences for non-English languages—up to 13 times longer than equivalent English texts—impacting inference cost and speed (Ahia et al., 2023; Sengupta et al., 2023; Petrov et al., 2023). Still, prior attempts to expand tokenizer vocabulary post-hoc are limited in number, and require substantial additional training (Kim et al., 2024; Zhao et al., 2024). In contrast, we propose to detect words the model “knows” and successfully detokenizes to a single vector, and use this to obtain new token embeddings.
Figure 6: Our 3-step method to expand LLM vocabulary without updates to core model parameters.
Our framework for vocabulary expansion follows a 3-step process (Fig. 6). Given a multi-token word $w$ that we would like to add to the vocabulary, and the model’s original input embedding and output unembedding matrices $E$ and $U$ with hidden dimension $d$, we (1) extract a detokenized, single-token representation $r$ for $w$ using a similar approach to Sec. 4.2: pass $w_{i}$ as input to the model and apply Patchscopes (Ghandeharioun et al., 2024) with prompt $P$ [^15] For $P$, we use the template “x x x x” where x is the patched representation of the new word. to the last token’s hidden states at all layers. We then identify $\ell$, the earliest layer at which the hidden state is successfully decoded into the full word, and set $r$ as that hidden state. Next, we (2.1) learn a set of linear maps $T_{\ell,E}$ and $T_{\ell,U}$ to project hidden states from layer $\ell$ of the model to the embedding and unembedding spaces, based only on the existing in-vocabulary tokens. [^16] We learn $T_{\ell,E}$ and $T_{\ell,U}$ by fitting an orthogonal procrustes transformation from the layer $\ell$ hidden states of all in-vocabulary tokens (when passed as inputs on their own), to their corresponding embeddings or unembeddings (see App. G for details). Importantly, these projections only rely on the model’s existing vocabulary, and do not depend on which multi-token words we choose to add to the vocabulary. Then, (2.2) for each detokenized representation $r$ taken from layer $\ell$, we apply $T_{\ell,E}$ and $T_{\ell,U}$ to obtain the initial entries in the embedding and unembedding matrices to represent the word, $\hat{e}$ and $\hat{u}$. Finally, after computing the initial entries for all new words, we (3) refine the new representations to obtain the final entries: we initialize two $d\times d$ all-zeros matrices $W_{E}$ and $W_{U}$, and set the new entries as $e=\hat{e}+W_{E}\hat{e}$ and $u=\hat{u}+W_{U}\hat{u}$. We train the refinement matrices $W_{E}$ and $W_{U}$ jointly in a short continued pretraining run, while keeping all other parameters frozen. Finally, we compute the final $e$ and $u$, and use these to represent the new word in the expanded vocabulary. Importantly, if a word is never successfully decoded from the hidden states in any of the layers in step (1), we assume it is not found in the model’s inner lexicon, and therefore do not add it to the vocabulary.
We apply our approach to Llama2-7B and experiment with three datasets: WikiText-103 (Merity et al., 2017), abstracts of biomedical articles from PubMed (Xiong et al., 2024), and the Arabic split of Wiki40B (Guo et al., 2020). For each dataset, we expand the model’s vocabulary with all multi-token words that appear at least $m$ times in the test set. [^17] We use $m=1,5,50$ for WikiText-103, PubMed and Arabic Wiki40B respectively. We then learn the refinement matrices $W_{E}$ and $W_{U}$ using 20M tokens from the train set. [^18] We use a sequence length of 512 and train on 10,000 sequences, taking up to 30 minutes on a single GPU. We evaluate models in a next-word prediction setup and compare our approach against two baselines: the original model with an unmodified vocabulary (original); and an expanded model that follows our framework but uses the mean embeddings of the word’s tokens in $E$ and $U$ to initialize its new representation, instead of the detokenized representations $r$, following Gee et al. (2022; mean embedding). For evaluation, perplexity is not a suitable metric, as differences in vocabulary between the approaches skew perplexity scores and prevent a fair comparison. Instead, we measure the model’s token-level top-1 accuracy when it is given each token’s previous context.
WikiText-103 PubMed Wiki40B-Arabic
Original Mean Embed. Ours Original Mean Embed. Ours Original Mean Embed. Ours
New words new token 0.071 0.171 0.123 0.180 0.117 0.402
original or new token 0.322 0.178 0.284 0.280 0.171 0.259 0.413 0.119 0.414
All words 0.522 0.473 0.519 0.517 0.479 0.511 0.535 0.211 0.532
Table 1: Token-level accuracy of finetuning-free vocabulary expansion for Llama2-7B on three datasets. We compare our method using detokenized representations (Ours) to the original model (Original) and to an expansion baseline using the word’s average token embeddings from $E$ and $U$ (Mean Embed.; Gee et al. 2022). Accuracy is reported for newly added words, where we distinguish between correctly predicting the new tokens (top row) and predicting either the word’s original first token, as the unexpanded model would, or its new token (middle row). We also report overall performance on all tokens (last row). Our method enables the frozen model to use the newly added input and output embeddings, while maintaining overall model performance.
Our results (Table 1) indicate that models can generalize to new vocabulary entries surprisingly well when these are initialized with their own detokenized representations. Unlike the mean embeddings expansion baseline, which both struggles with using new tokens and degrades overall performance, our method allows the model to successfully integrate new vocabulary words while preserving its accuracy on existing words. This effect is particularly pronounced in Arabic Wiki40B, where the model almost always selects new tokens instead of the original tokenization (40.2% vs 41.3%), demonstrating the potential of our approach for multilingual and domain-specific applications.
Our method also shows potential to reduce the number of tokens processed during encoding and inference: across the three datasets, we observe a reduction in the total number of tokens processed during encoding by 10.5% to 14.5% (see Table 5 in App. H), while maintaining model performance. In future work, we will further explore the application of our framework to multilingual adaptation (Petrov et al., 2023; Alabi et al., 2022) and continual domain-adaptive pretraining (Ke et al., 2023; Yıldız et al., 2024; Gururangan et al., 2020).